# Xuezhang Chen  and Shihong Zhang

### v1

@@ 3.2 Noncritical dimension: The unit ball
Clearly, $F$ is exact an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and thus $F=F^{-1}$.
The following important but elementary identity will be frequently used:
@@
This can be verified by a straightforward calculation.
Given $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, we set
$u=U\circ FU_{0}.$ (3.9)
Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $v$ belong to $O(|X|^{3-n-k})$ for every $k\in\mathbb{N}$. Moreover, by conformal covariance of the Paneitz operator, we have
$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U\circ FU_{0})=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
@@
$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad 0\leq k\leq 3.$ (3.10)
Proof of Theorem 1.2. For assertion $(i)$, we first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
@@
$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F\left((x,0)\right)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F((x,0)).$
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=f_{0}\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=f_{2}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.11)
As a direct consequence of Remark 2.2, the solution to the above problem (3.11) is unique. This together with Theorem 1.1 gives
@@
$\eta=$ $F((y,0))=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
@@
and from (3.8) that
@@
As indicated in the preceding discussion, for other pairs of conformal boundary operators it suffices to show the uniqueness.
The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ is trivial. For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, the uniqueness follows from a similar argument above combined with Lemma 2.3 or Remark 2.1.
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}^{2}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
@@
So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well.
For assertion $(ii)$, the multiplicity of smooth solutions for the remaining two boundary conditions is a direct consequence of the nex Lemma 3.2. $\Box$
@@
The following homogeneous biharmonic boundary value problems:
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.12)
@@
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
@@
As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\partial_{t}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
@@
This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, under which boundary conditions problem (3.12) or (3.13) fails to have the uniqueness of the solution as explained in [1].
3.3 Critical dimension
Motivated by Chen-Shi [10, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
Proof of Theorem 1.4. Based on the proof of Theorems 1.1 and 1.2, it suffices to point out several significant differences.
(1) Observe that
to show
(2) Suppose $U$ is a solution of (1.7). In critical dimension four, we introduce
@@ 3.3 Critical dimension
$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
This essentially follows from conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ that
@@
$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}).$
In particular, $(T_{1}^{3})_{g_{\mathbb{S}^{3}}}=1,(T_{2}^{3})_{g_{\mathbb{S}^{3}}}=2,(T_{3}^{3})_{g_{\mathbb{S}^{3}}}=4$.
Suppose $u$ is a solution of (1.6). By the biharmonic Poisson kernel on $\mathbb{R}_{+}^{4}$ we have
@@
$I(x,t)=\begin{cases}\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{g_{\mathbb{S}^{3}}}]\mathrm{d}y,\\ \int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y\quad\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
@@
Hence, using variable changes as in the proof of Theorem 1.2 we obtain
@@
It remains to determine the constant $C$ in the following way: On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [10, Theorem 1(1)], more explicitly,
@@
Integrating the above equation over $\mathbb{S}^{3}$ to give
@@
where we have used (3.15) and the fact that $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi),\forall~{}\eta\in\mathbb{S}^{3}$ is constant. This follows that
@@
Then the desired assertion follows by inserting the above constant into (3.14). $\Box$
It is not hard to see that Theorem 1.4 could give a direct proof of [6, Proposition 2.2, Lemma 2.3]).
We point out that the proof of Theorems 1.1 and 1.2 is still valid for the case of $n=2$. However, we shall not go deep into this, since it has been well studied by authors [11].
Thanks to biharmonic Poisson kernels on $\mathbb{R}^{n+1}$, we can recover a special case in [7, Theorem 1.4], which clarifies the relationship between extrinsic and intrinsic GJMS operators, and leave its proof to Appendix A for clarity. Something about $P_{1}$ and $P_{3}$.
Proposition 3.1 (cf. [7, Theorem 1.4] and [11, Porposition 2.2]).
@@
where $P_{1},P_{3}$ are the fractional GJMS operators with respect to the round metric, see [7, 10].
Proposition 3.1 indicates that a Dirichlet datum is equivalent to a $\mathscr{B}_{3}^{3}$ datum since $P_{3}$ is invertible; a $\mathscr{B}_{1}^{3}$ datum is equivalent to a $\mathscr{B}_{2}^{3}$ datum since $P_{1}$ is invertible. That is exact the reason why we can construct multiple smooth solutions for a biharmonic equation on $\mathbb{B}^{n+1}$ with these two types of boundary conditions.

### v2

@@ 3.2 The unit ball
Clearly, $F$ is an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and its inverse $F^{-1}=F$.
The following elementary identity will be frequently used:
@@
This follows by a straightforward calculation.
For $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ we let
$u=U_{0}(U\circ F).$ (3.9)
Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $u$ belong to $O(|X|^{3-n-k})$ for $k\in\mathbb{N}$. Moreover, by conformal covariance of the Paneitz operator, we have
$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U_{0}U\circ F)=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
@@
$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad k\in\{0,1,2,3\}.$ (3.10)
Theorem 3.1.
Suppose $n\geq 2$ and $n\neq 3$, and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
Then for $i,j\in\{0,1,2,3\},i<j,i+j\neq 3$, the solution of the following BVP
$\begin{cases}\displaystyle\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}&\mathrm{~{}~{}on~{}~{}}\quad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\mathbb{S}^{n},\end{cases}$ (3.11)
is unique and has the form of
Proof..
We first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
@@
$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F(x,0)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F(x,0).$
@@
$\begin{cases}\displaystyle\qquad~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\qquad\quad~{}~{}u=f_{0}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u-\overline{\Delta}u=f_{2}\qquad\quad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.12)
As a direct consequence of Remark 2.2, the above problem (3.12) admits a unique solution. This together with Theorem 1.1 gives
@@
$\eta=$ $F(y,0)=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
@@
and
$|\xi+\mathbf{e}_{n+1}|^{2}=\frac{4}{(1+t)^{2}+|x|^{2}},\quad|\eta+\mathbf{e}_{n+1}|^{2}=\frac{4}{1+|y|^{2}}.$
Moreover, it follows from (3.8) that
@@
As indicated in the preceding discussion, for other proper conformal boundary operator pairs it remains to show the uniqueness.
The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ directly follows by an integration by parts. For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, the uniqueness follows from a similar argument above combined with Lemma 2.3 or Remark 2.1.
@@
$\begin{cases}\displaystyle\qquad\qquad~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\qquad\partial_{t}^{2}u-\overline{\Delta}u=0&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u=0\qquad\quad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\end{cases}$
@@
So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well. $\Box$
Remark 3.2.
It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function for the fractional GJMS operator $P_{3}$ of order three on $\mathbb{S}^{n}$ (cf. the first author and Shi [13, Theorem 1(3)]).
@@
The following homogeneous biharmonic BVPs:
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad\quad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad\quad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.14)
@@
As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ satisfies both
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\quad~{}u=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle~{}~{}\partial_{t}u=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u-\overline{\Delta}u=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
@@
This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$. Due to this reason, problem (3.13) or (3.14) fails to have the uniqueness.
Motivated by Shi and the first author [13, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
Theorem 3.2.
Suppose $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{3})$ for $i\in\{0,1,2,3\}$. Then the biharmonic Poisson kernels $\bar{P}_{i}^{3}$ for $\mathbb{B}^{4}$ have the same representation formulae as in Theorem 3.1 except for
where $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
  1. (i)
    $i=1,2$ and $j=3$, the solution of (3.11) is
  2. (ii)
    $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (3.11) is
Proof..
Suppose $U$ is a smooth solution of (3.11). For $n=3$, we introduce
@@
$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{\mathbb{S}^{3}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
By the conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ we have
@@
$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{\mathbb{S}^{3}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{\mathbb{S}^{3}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{\mathbb{S}^{3}}).$
In particular, $(T_{1}^{3})_{\mathbb{S}^{3}}=1,(T_{2}^{3})_{\mathbb{S}^{3}}=2,(T_{3}^{3})_{\mathbb{S}^{3}}=4$.
Suppose $u$ is a solution of (1.6). Using the biharmonic Poisson kernel for $\mathbb{R}_{+}^{4}$ we have
@@
$I(x,t)=\begin{cases}\displaystyle\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{\mathbb{S}^{3}}]\mathrm{d}y,\\ \displaystyle\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y~{}~{}\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
@@
Hence, using the variable changes as in the proof of Theorem 3.1 we obtain
@@
It remains to determine the constant $C$. On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [13, Theorem 1(1)], more explicitly,
@@
Integrating the above equation over $\mathbb{S}^{3}$ gives
@@
where we have used (3.16) and the fact that $\forall~{}\eta\in\mathbb{S}^{3}$, $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi)$ is constant. This follows that
@@
Then the desired assertion follows by substituting the above constant into (3.15). $\Box$
Finally, we point out that the proof of Theorems 1.1 and 3.1 is still valid for $n=2$. However, we shall not go deep into this, since it has been well studied by the authors [15].
For a Poincaré-Einstein manfold $(X^{n+1},g_{+})$ with conformal infinity $(\partial X=M,[h])$, and assume $(n^{2}-1)/4,(n^{2}-9)/4\notin\sigma_{\mathrm{pp}}(-\Delta_{g_{+}})$, with the factorization of $P_{4}^{g_{+}}$ and explicit formulae of extrinsic GJMS boundary operators $\mathscr{B}_{k}^{3}$, Case [9] used the scattering operator to find fractional GJMS operators $P_{1}^{g},P_{3}^{g}$ from directly computing $(\mathscr{B}_{k}^{3})_{g}u$, where $P_{4}^{g}u=0$ in $X$ for $g=r^{2}g_{+}$ with $r$ being a geodesic defining function. As for the Poincaré ball $(\mathbb{B}^{n+1},g_{+}=(\frac{2}{1-|x|^{2}})^{2}|\mathrm{d}x|^{2})$, the spectrum $\sigma(-\Delta_{g_{+}})=[n^{2}/4,+\infty)$ and [9, Theorem 1.4] is applicable.
Proposition 3.1.
@@
where $P_{1},P_{3}$ are the first- and third-order GJMS operators with respect to the round metric, see [9, 13].
An alternative proof of Proposition 3.1 using the biharmonic Poisson kernel is available in Appendix A.